Aug 18th, 09:50: Systems are working normally, monitoring showed no issues in the past 24 hours; marking incident as fixed.
Aug 17th, 10:02: High incoming traffic has been mitigated, services should be operating normally again. We are monitoring, but the issue seems to be fixed.
Aug 17th, 09:43: A fix has been applied, BGP seems to be working, but we are seeing extreme amounts of traffic, and are verifying.
Aug 17th, 09:30: The problem is related to BGP and the new routers. We are in the process of applying a fix.
Aug 17th, 09:18: There are network problems which seem to be caused by the BGP routing table. We are investigating.